# Persistent Homology for Structural Characterization in Disordered Systems

### v3

@@ III.1.1 ML Results and Analysis
The PCA significantly reduces the number of predictors with minimal performance loss, which suggests that particle rearrangement (Task 1) and phase (Task 2) are determined by a small number of features. With $p=98\%$, the number of features was reduced to 188 in Task 1 and 122 in Task 2, while maintaining high accuracy—85% for Task 1 and 95% for Task 2.
III.1.2 Feature Importance Analysis (FIA)
We use the Shapley values introduced in Sec. II.4.4 to evaluate the importance of different homology classes, specifically $H_{0}$, $H_{1}$ and $H_{2}$, in two classification tasks.
The results listed in Table 2 indicate that the significance of $H_{0}$ is slightly lower than that of $H_{1}$ and $H_{2}$, with $H_{1}$ and $H_{2}$ together contributing over 70% of the total importance in both tasks.
@@ III.1.2 Feature Importance Analysis (FIA)
Our further experiments indicated that in Task 1, at least 7 predictors were needed to maintain classification accuracy, while in Task 2, only a single predictor was sufficient to achieve an accuracy of 94.2%. This suggests that three-phase classification of multi-particle system is primarily driven by this one critical predictor.
This predictor, corresponding to the largest eigenvalue from the PCA, is the 1389th element of $\mathbb{I}_{\text{entire}}^{(P)}$, located at pixel (28, 34) in the PI. Given the PI dimensions of $40\times 40$ and birth and persistence ranges of $[-0.1,1.5]$, each pixel has a resolution of $0.04$. We can get that the pixel $(28,34)$ is with birth range of $[1.02,1.06]$ and persistence ranges of and $[0.1,0.14]$ in the PD, respectively. The feature importance, based on the eigenvalues of the principal components, revealed that this key predictor corresponds to the 1389th column of the original feature matrix, which is the 1389th element of the vector $\mathbb{I}_{\text{entire}}^{(P)}$.
III.2 Non-ML Approach
@@
In Sec. III.1.2, we identified a single variable that enables highly accurate phase classification (accuracy of $94.2\%$), showing that this variable alone can act as a key factor in distinguishing phases. However, the mapping trained using ML is not explicit, making it difficult to express its analytical form.
The SI can be seen as an analytical substitute for the mapping trained by ML. While its accuracy is slightly lower, it can still achieve near-perfect classification between liquids and solids. Additionally, the overlapping regions between amorphous and crystalline phases lie outside the interquartile range, indicating that SI can, in most cases, achieve reasonably accurate three-phase classification. Although SI sacrifices some accuracy, it significantly enhances interpretability.
@@
III.3 Metrics on Trajectories
@@
We tracked the evolution in four metrics including normalized MSD, $Q_{6}(t)/\mathscr{N}$, $\text{SI}(t)$ and $\mathcal{S}(t)$ over time on ten trajectories in test group (Group 2), i.e., Traj($\text{Traj}(T^{(2)}_{k}),k\in[1,10]\cap\mathbb{N}$, as well as two trajectories in Group 3, i.e., $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$. The data for all metrics are provided in Fig. 17 of Appendix C for reference.
The trajectories in the test set (Group 2) are divided into two categories: Category I, with $k\leq 6$, undergoes complete crystallization from an amorphous or liquid state, while Category II, with $k\geq 7$, remains in the liquid phase. In Fig. 9, the blue solid line shows the average of SI (Fig. 9a)) and Global Softness (Fig. 9b)) for each trajectory in Category I, while the orange solid line shows the same for Category II. In Category I, SI rises sharply as Global Softness decreases, whereas in Category II, SI stays low and stable, with Global Softness remaining at a high positive level. This distinction clearly differentiates crystallized states from non-crystallized ones. Notably, the Global Softness was calculated using a linear SVM explaining 98% of the explained variance, trained in Task 1. In contrast, other three metrics are non-ML based.
In Fig. 8, without loss of generality, we report the evolution of four metrics over time for four representative trajectories: (i) $\text{Traj}(T_{3}^{(2)})$, which undergoes a complete transition from amorphous to crystalline (Fig. 8a); (ii) $\text{Traj}(T_{6}^{(2)})$, which undergoes a complete crystallization from liquid (Fig. 8b); (iii) $\text{Traj}_{\text{KA}}^{(1)}$, depicting a glass transition from a liquid (Fig. 8c); and (iv) $\text{Traj}_{\text{KA}}^{(2)}$, which remains in the liquid phase during quenching (Fig. 8d).
@@ A.2 The Parameter Selection for Rearrangement Tendency (Task 1)
Here, the parameters we need to set are $t_{R/2}$ and $p_{c}$, aligned with Ref. [80]. Fig. 1314, and 15 are all adapted from Ref. [80]. The settings is based on the statistics on Group 1 (Set 1), and we cannot use the data in the Group 2 (Set 2) at this stage to avoid the data leakage.
@@ Appendix B Correctness of Global Softness
Appendix C Extra Data

### v4

@@ IV.1.2 ML Results and Analysis
As shown in Table 2, further reducing the explained variance ratio $z$ reveals that at least 7 predictors are required to prevent a significant decline in the accuracy of Task 1. Additionally, we evaluated scenarios with the number of predictors ranging from 1 to 7, with detailed results provided in Table 6 of Appendix E. Our analysis indicates that at least 3 predictors are necessary for the accuracy of Task 1 to be significantly higher than that of random selection, confirming that a single predictor is insufficient. This suggests that mapping particles to a binary classification (soft or hard) using a single variable is not feasible.
In contrast, Task 2 can be effectively performed using a single variable without a significant loss in accuracy. This suggests that mapping the entire system to a three-phase classification (crystal, amorphous, or liquid) based on a single variable is feasible. However, deriving such a mapping analytically remains still challenging.
IV.1.3 Feature Importance Analysis (FIA)
Here, we address two key questions: first, the significance of homology classes of different dimensions in the two classification tasks; and second, in the three-phase classification task of Task 2, which column in the feature matrix plays a decisive role—specifically, which variable serves as the primary determining factor.
Firstly, we use the Shapley values introduced in Sec. II.3.5 to evaluate the importance of different homology classes, specifically $H_{0}$, $H_{1}$ and $H_{2}$, in two classification tasks. The results listed in Table 3 indicate that the significance of $H_{0}$ is slightly lower than that of $H_{1}$ and $H_{2}$, with $H_{1}$ and $H_{2}$ together contributing over 70% of the total importance in both tasks.
@@ IV.1.3 Feature Importance Analysis (FIA)
Secondly, in Task 2, a single predictor is sufficient to achieve an accuracy of 94.2%, suggesting that the three-phase classification of a multi-particle system is largely dictated by this single critical predictor. We conclude that this variable corresponds to the 1389th element of the vector $\mathbb{I}_{\text{entire}}^{(P)}$, with the reasoning process detailed in Appendix C. This single variable, in conjunction with the SVM mapping $f$, effectively characterizes the global phase structure.
IV.2 Descriptive Statistics on PH (non-ML approach)
In Sec. II.4, we introduced a mapping from the results of the PH analysis to a single real number in the range $[0,+\infty)$, specifically the Separation Index (SI), as defined in Eq.9. Since both the neighborhood of a particle and the entire system can be represented as point clouds, the SI applies to both local and global structures, mapping each to a single real number.
@@ IV.2 Descriptive Statistics on PH (non-ML approach)
The SI can be seen as an analytical approximation for the mapping trained by ML. For Task 2 (three-phase classification), while its accuracy is slightly lower, it can still achieve near-perfect classification between liquids and solids (Fig. 5b)). Additionally, the overlapping regions between amorphous and crystalline phases lie outside the interquartile range (Fig. 5b)), indicating that SI can, in most cases, achieve reasonably accurate three-phase classification. Although SI sacrifices some accuracy, it significantly enhances interpretability. For Task 1, although we have rigorously demonstrated that a single variable is insufficient for fully classifying particles on {soft, hard}, the interquartile ranges of the SI for soft and hard particles begin to separate as the neighborhood radius increases (Fig.5a)). Simultaneously, their distribution patterns progressively resemble those of liquids and solids (compare Fig.5a) to Fig. 5b)).
Therefore, beyond providing an approximate analytical mapping for classification, SI plays a more crucial role as a bridge, enabling the analysis of the relationship between the soft-hard property of a particle and the global phase structure within a unified framework. This analysis is visualized in Fig. 5.
Although hard particles tend to be located in symmetric environments, the tendency for rearrangement and the preservation of order are fundamentally distinct properties. Although symmetry is related to particle rigidity, other factors—such as local packing density and stress distribution—can also significantly influence rearrangement behavior. Therefore, $q_{6}$, which primarily measures six-fold symmetry, is not a reliable predictor of particle rearrangements, as demonstrated in Appendix D.
@@
IV.3 Metrics on Trajectories
Tracking the evolution of various metrics along molecular dynamics trajectories is a widely used approach for studying how target properties evolve over time under controlled conditions. In this study, we leverage this method to identify the onset of phase transitions using our proposed metrics, such as SI and Global Softness, while also providing a comparison with traditional metrics like $Q_{6}$ and MSD.
@@ IV.3 Metrics on Trajectories
We tracked the evolution in four metrics including normalized MSD, $Q_{6}(t)/\mathscr{N}$, $\text{SI}(t)$ and $\mathcal{S}(t)$ over time on ten trajectories in test group (Group 2), i.e., Traj($\text{Traj}(T^{(2)}_{k}),k\in[1,10]\cap\mathbb{N}$, as well as two trajectories in Group 3, i.e., $\text{Traj}_{\text{KA}}^{(1)}$ and $\text{Traj}_{\text{KA}}^{(2)}$. The data for all metrics are provided in Fig. 19 of Appendix E for reference.
The trajectories corresponding to the test set of Group 2 are divided into two categories: Category I, with $k\leq 6$, undergoes complete crystallization from an amorphous or liquid state, while Category II, with $k\geq 7$, remains in the liquid phase. In Fig. 9, the blue solid line shows the average of SI (Fig. 9a)) and Global Softness (Fig. 9b)) for each trajectory in Category I, while the orange solid line shows the same for Category II. In Category I, SI rises sharply as Global Softness decreases, whereas in Category II, SI stays low and stable, with Global Softness remaining at a high positive level. This distinction clearly differentiates crystallized states from non-crystallized ones. Notably, the Global Softness was calculated using a linear SVM explaining 98% of the explained variance, trained in Task 1. In contrast, other three metrics are non-ML based.
In Fig. 10, without loss of generality, we report the evolution of four metrics over time for four representative trajectories: (i) $\text{Traj}(T_{3}^{(2)})$, which undergoes a complete transition from amorphous to crystalline (Fig. 10a)); (ii) $\text{Traj}(T_{6}^{(2)})$, which undergoes a complete crystallization from liquid (Fig. 10b)); (iii) $\text{Traj}_{\text{KA}}^{(1)}$, depicting a glass transition from a liquid (Fig. 10c)); and (iv) $\text{Traj}_{\text{KA}}^{(2)}$, which remains in the liquid phase during quenching (Fig. 10d)).
@@ A.2 The Parameter Selection for Rearrangement Tendency (Task 1)
Here, the parameters we need to set are $t_{R/2}$ and $p_{c}$, aligned with Ref. [94]. Fig. 1415, and 16 are all adapted from Ref. [94]. The settings is based on the statistics on Group 1 (Set 1), and we cannot use the data in the Group 2 (Set 2) at this stage to avoid the data leakage.
@@ Appendix B Correctness of Global Softness
Appendix C Details of Feature Importance Analysis (FIA)
The single decisive predictor mentioned in Sec. IV.1.3, corresponding to the largest eigenvalue from the PCA, is the 1389th element of $\mathbb{I}_{\text{entire}}^{(P)}$, located at pixel (28, 34) in the PI. Given the PI dimensions of $40\times 40$ and birth and persistence ranges of $[-0.1,1.5]$, each pixel has a resolution of $0.04$. We can get that the pixel $(28,34)$ is with birth range of $[1.02,1.06]$ and persistence ranges of and $[0.1,0.14]$ in the PD, respectively. The feature importance, based on the eigenvalues of the principal components, revealed that this key predictor corresponds to the 1389th column of the original feature matrix, which is the 1389th element of the vector $\mathbb{I}_{\text{entire}}^{(P)}$.
Appendix D comparative Analysis Based on the $q_{6}$-SI Swapped State Sets
[…]
Figure 18: Distribution of $q_{6}$ over the state set {soft, hard} and the Separation Index (SI) over the state set {ordered, disordered}. Note that the dataset used here is the same as in the Fig. 5a), with only new metrics and labels applied.
As shown in Fig. 18a), aside from the slightly lower average $q_{6}$ value observed for soft particles compared to hard particles, no other significant non-trivial patterns are evident. Thus, $q_{6}$ is not an effective predictor of particle rearrangement tendencies, nor does it serve as a reliable criterion for distinguishing between soft and hard particles. While the results suggest that hard particles exhibit, on average, a higher degree of local symmetry, this difference alone is insufficient to justify its use as a classification metric.
In Fig. 18b), although the average SI differs between ordered and disordered particles, the lower bound of the IQR for ordered particles does not separate significantly from the upper bound of the IQR for disordered particles as the neighborhood radius increases. Instead, it does not exhibit any clear non-trivial trends. Therefore, while particles in ordered environments generally display higher average SI values, this distinction also falls short of providing a robust basis for classification.
The particle rearrangement tendency and the degree of six-fold symmetry preservation in the local environment represent entirely different properties and cannot be used interchangeably as classification criteria. Overall, SI exhibits slightly better performance than $q_{6}$.
Appendix E Extra Data
Table 5 provides the values of the 20 monotonically increasing temperatures mentioned in Sec. III.2.
$T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ $T_{7}$ $T_{8}$ $T_{9}$ $T_{10}$
0.50 0.53 0.55 0.57 0.60 0.63 0.65 0.67 0.70 0.73
$T_{11}$ $T_{12}$ $T_{13}$ $T_{14}$ $T_{15}$ $T_{16}$ $T_{17}$ $T_{18}$ $T_{19}$ $T_{20}$
0.75 0.77 0.80 0.83 0.85 0.88 0.90 0.93 0.95 0.97
Table 5: The 20 different temperatures used in Protocol 1 for generating LJ systems.
Table 6: Classification performance of RBF kernel SVM after PCA with different explained variances in Task 1.
VAR Explained Num of Predictors Accuracy MCC
98% 188 84.2% 0.686
50% 7 82.4% 0.649
48% 6 81.3% 0.626
45% 5 80.7% 0.614
40% 4 80.3% 0.607
35% 3 79.5% 0.589
30% 2 51.2% 0.030
23% 1 50.3% 0.007
As demonstrated in Table. 6, in Task 1, after applying PCA with different variance explained ratios to the feature matrix, we assessed the classification performance of an SVM with an RBF kernel. The results indicate that at least 3 variables are required to characterize the hard-or-soft property of the particles with an accuracy significantly higher than random selection. Furthermore, to achieve an accuracy loss within 2% compared to the baseline (the first row of Table 6), at least 7 variables are necessary.